OpenStudy (anonymous):

Find m<1.m<2,m<3,m<4 Where m<2 is equal to (x+2) and m<3 is equal to (3x+18)

OpenStudy (diyadiya):

DO you have a figure? If yes, Please attach it!

OpenStudy (anonymous):

there you go pls answer

OpenStudy (apoorvk):

@Faye281 vertically opposite angles are equal in measure, and adjacent angles add up 180 degrees in a system of intersecting lines - know this?

OpenStudy (anonymous):

yeah

OpenStudy (diyadiya):

Vertically opposite sides are equal (m<1 = m<3 ) &( m<2 = m<4) and m<1 + m<2 = 180 m<2 + m<3 = 180 m<3 + m<4 = 180 m<4 + m<1 = 180 m<2 + m<3 = 180 x + 2 + 3x + 18 = 180 can you find x now?

OpenStudy (apoorvk):

 for sake of ease.